OpenStudy (anonymous):

A bag contains 5 red marbles and 8 yellow marbles. You are asked to draw 3 marbles from the bag without replacement. In how many ways can you draw 2 reds and 1 yellow?

OpenStudy (anonymous):

how many ways, or probability?

OpenStudy (anonymous):

how many ways: you have 5 red marbles and so there are \(_5C_2=10\) ways of selecting two of them then there are \(_8C_1=8\) ways of selecting the one yellow one out of the 8 multiply to get your answer

OpenStudy (anonymous):

i think @Nurali made a typo, and meant \(_8C_1=8\) as opposed to \(_6C_1=6\) for the yellow one

OpenStudy (nurali):

5C2*8C1 = 5*4/(2*1)*8 = 80 ways.

OpenStudy (anonymous):

its how many ways. possible answers: 160 40 17 4,838,400 200

OpenStudy (magbak):

This question Is not a valid question with correctly valid answers

OpenStudy (anonymous):

maybe we are to assume that order counts in this problem, in which case the answer would be different

OpenStudy (magbak):

Even if order is to matter because that is the way I presumed it there will only be 3 ways

OpenStudy (anonymous):

kind of strange though, since it is written in such a way that would make you think order does not count, as you cannot tell the two reds apart

OpenStudy (anonymous):

Basically this is why I'm confused. I didn't know if it was just me or...

OpenStudy (anonymous):

but if order counts we can do that too

OpenStudy (magbak):

That is what I did and resulted with only 3 possible ways if both reds are the same.

OpenStudy (magbak):

And if it was two different reds then their is still only 6 possible way to draw the marbles for the bag.

OpenStudy (anonymous):

but i don't get any of the above answers if order counts it could be red red yellow red yellow red yellow red red for each the number of ways would be \(5\times 4\times 8=160\)

OpenStudy (anonymous):

so in this case the answer should be (i think) \(3\times 160=480\)

OpenStudy (magbak):

Actualy 160 is a answer choice

OpenStudy (anonymous):

yeah, but that is not the answer if order counts

OpenStudy (magbak):

Yes true but this is probably a pulication error or a mistake from the her part

OpenStudy (anonymous):

if order counts, then red red yellow is different from red yellow red is different from yellow red red either order counts, or it doesn't if it doesn't the answer is \(80\)